# U-DiTs: Downsample Tokens in U-Shaped Diffusion Transformers

### v2

@@ 1 Introduction
Interestingly, DiTs have discarded the U-Net architecture [27] that is universally applied in manifold previous works, either in pixel [18; 12] or latent space [26]. The use of isotropic architectures in DiTs is indeed successful, as scaled-up DiT models achieve supreme performance. However, the abandonment of the widely-applied U-Net architecture by DiTs and their improvements [17; 9; 23] on latent-space image generation tasks triggers our curiosity, because the U-Net inductive bias is always believed to help denoising. Hence, we rethink deploying DiTs on a canonical U-Net architecture.
@@
Hence, we rethink the self-attention mechanism in DiT-UNet. The backbone in a latent U-Net denoiser provides a feature where low-frequency components dominate [28]. The discovery implies the existence of redundancies in backbone features: the attention module in the U-Net diffuser should highlight low-frequency domains. As previous theories praised downsampling for filtering high-frequency noises in diffusion [36], we seek to leverage this natural low-pass filter by performing token downsampling on the features for self-attention. Unlike previous transformer works [16; 39; 29] that downsample key-value pairs only, we radically downsample the query-key-value tuple altogether, such that self-attention is performed among downsampled latent tokens. It is surprising that when we incorporate self-attention with downsampled tokens into DiT-UNet, better results are achieved on latent U-Net diffusers with a significant reduction of computation.
@@ 2 Preliminaries
Vision Transformers. ViTs [14] have introduced a transformer backbone to vision tasks by patchifying the input and viewing an image as a sequence of patch tokens and have proved its effectiveness on large-scale image classification tasks. While ViTs adopt an isotropic architecture, some following works on vision transformers [34; 22] propose a pyramid-like hierarchical architecture that gradually downsamples the feature. The pyramid architecture has proved highly effective in classification and other downstream tasks.
Vision transformers are also mainstream backbones for denoising models. IPT [6] introduces an isotropic transformer backbone for denoising and other low-level tasks. Some later works [20; 19; 8] follow the isotropic convention, but other denoising works [35; 37] shift to U-Net backbones as their design. The pioneering work of U-ViT [1] and DiT [25] introduces full-transformer backbones to diffusion as denoisers.
Recent Advancements in Diffusion Transformers. Following DiTs, some works investigate the training and diffusion [15; 24] strategies of Diffusion Transformers. Other works focus on the design of the DiT backbone. DiffiT [7; 17] introduces a new fusion method for conditions; FiT [23] and VisionLLaMA [9] strengthens DiT by introducing LLM tricks including RoPE2D [31] and SwishGLU. These transformer-based diffusion works agree on adopting isotropic architectures on latents, i.e. the latent feature space is not downsampled throughout the whole diffusion model. The authors of DiT [25] even regard the inductive bias of U-Net as “not crucial”.
U-Nets for Diffusion. From canonical works [18; 30; 12; 26], the design philosophy of U-Net [27] is generally accepted in diffusion. Specifically, Stable Diffusion [26] uses a U-Net-based denoiser on the compressed latent space for high-resolution image synthesis, which is highly successful in manifold generative tasks. Some previous trials on diffusion transformers [4; 17; 10] also adopt U-Net on pixel-space generation tasks; but strangely, they shifted to isotropic DiT-like structures for latent-space diffusion. Despite its popularity in pixel-space diffusion, the U-Net architecture is not widely accepted in recent transformer-oriented works on latent-space diffusion.
Motivated by this, we are dedicated to investigating the potential of Transformer-backboned U-Net on latent-space diffusion.
@@ 3 Investigating U-Net DiTs in Latent
In fact, attention to downsampled tokens is not new. Previous works regarding vision transformers [16; 39] have proposed methods to downsample key-value pairs for computation cost reduction. Recent work on training-free acceleration of diffusion [29] also applies key-value downsampling on Stable Diffusion models. But these works maintain the number of queries, and thus the downsampling operation is not completely performed. Besides, these downsampling measures usually involves a reduction of tensor size, which could result in a significant loss in information.
@@
Complexity analysis. Apart from the performance benefits, we are aware that downsampled self-attention could save as much as 1/3 of the overall computation cost compared to full-scale self-attention. We conduct a brief computation complexity analysis on the self-attention mechanism to explain where the savings come from.
@@
In the proposed self-attention on downsampled tokens, four sets of downsampled query-key-value tuples $4\times\left(Q_{\downarrow 2},K_{\downarrow 2},V_{\downarrow 2}\right)\in\mathbb{R}^{(\frac{N}{2})^{2}\times d}$ performs self-attention respectively. While each self-attention operation costs only 1/16 of full-scale self-attention, the total cost for downsampled self-attention is 1/4 of full-scale self-attention. 3/4 of the computation costs by self-attention is saved via token downsampling.
@@ 4 Scaling the Model Up
Settings. We adopt the training setting of DiT. The same VAE (i.e. sd-vae-ft-ema) for latent diffusion models [26] and the AdamW optimizer is adopted. The training hyperparameters are kept unchanged, including global batch size 256, learning rate $41e-$ 4, weight decay 0, and global seed 0. The training is conducted with the training set of ImageNet 2012 [11].
Apart from the self-attention on downsampling as introduced in the toy experiment (Section 3), we further introduce a series of modifications to U-DiTs, including cosine similarity attention [21; 19], RoPE2D [31; 23; 9], depthwise conv FFN [35; 3; 39], and re-parametrization [13; 32]. The contribution of each modification is quantitatively evaluated in Sec. 8.
@@ 4.1 U-DiT at Larger Scales
Table 2: Comparing U-DiTs against DiTs on ImageNet 256$\times$256 generation. Experiments with a supermark are replicated according to the official code of DiT. We compare models trained for 400K iterations with the standard training hyperparameters of DiT. The performance of U-DiTs is outstanding: U-DiT-B could beat DiT-XL/2 with only 1/6 of inference FLOPs; U-DiT-L could outcompete DiT-XL/2 by $10$ FIDs.
@@
It could be concluded from Tab. 2 that all U-DiT models could outcompete their isotropic counterparts by considerable margins. Specifically, U-DiT-S and U-DiT-B could outperform DiTs of comparable size by $\sim 30$ FIDs; U-DiT-L could outperform DiT-XL/2 by $\sim 10$ FIDs. It is shocking that U-DiT-B could outcompete DiT-XL/2 with only 1/6 of the computation costs. To present the advantage of our method better, we also include the performance of U-DiTs in an FID-50K versus FLOPs plot (Fig. 2). Apart from DiTs and U-DiTs, we also include other state-of-the-art methods: SiT [24] that proposes an interpolant framework for DiTs, and SiT-LLaMA [9] that combines state-of-the-art DiT backbone VisionLLaMA and SiT. The advantages of U-DiTs over other baselines are prominent in the plot. The results highlight the extraordinary scalability of the proposed U-DiT models.
@@ 4.2 Ablations
Next, we inspect other modifications that further refine U-DiTs (lower part of Tab. 8). Swin Transformer V2 [21] proposes a stronger variant of self-attention: instead of directly multiplying Q and K matrices, cosine similarities between queries and keys are used. We apply the design to our self-attention, which yields $\sim 2.5$FIDs of improvement. RoPE [31] is a powerful positional embedding method, which has been widely applied in Large Language Models. Following the latest diffusion transformer works [23; 9], we inject 2-dimensional RoPE (RoPE2D) into queries and keys right before self-attention. The introduction of RoPE2D improves performance by $\sim 2.5$FIDs. Some recent transformer works strengthen MLP by inserting a depthwise convolution layer between two linear mappings [35; 3; 39]. As the measure is proved effective in these works, we borrow it to our U-DiT model, improving $\sim 5$FIDs. As re-parametrization during training [13] could improve model performance, we apply the trick to FFN [32] and bring an additional improvement of $\sim 3.5$FIDs. Above all, based on these components, the proposed U-DiTs are further improved.

### v3

@@ 1 Introduction
Interestingly, DiTs have discarded the U-Net architecture [30] that is universally applied in manifold previous works, either in pixel [20; 13] or latent space [29]. The use of isotropic (i.e. standard transformer; a plain stack of transformer blocks) architectures in DiTs is indeed successful, as scaled-up DiT models achieve supreme performance. However, the abandonment of the widely-applied U-Net architecture by DiTs and their improvements [18; 10; 26] on latent-space image generation tasks triggers our curiosity, because the U-Net inductive bias is always believed to help denoising. Hence, we rethink deploying DiTs on a canonical U-Net architecture.
@@
Hence, we rethink the self-attention mechanism in DiT-UNet. The backbone in a latent U-Net denoiser provides a feature where low-frequency components dominate [31]. The discovery implies the existence of redundancies in backbone features: the attention module in the U-Net diffuser should highlight low-frequency domains. As previous theories praised downsampling for filtering high-frequency noises in diffusion [39], we seek to leverage this natural low-pass filter by performing token downsampling on the features for self-attention. Unlike previous transformer works [17; 44; 32] that downsample key-value pairs only, we radically downsample the query-key-value tuple altogether, such that self-attention is performed among downsampled latent tokens. It is surprising that when we incorporate self-attention with downsampled tokens into DiT-UNet, better results are achieved on latent U-Net diffusers with a significant reduction of computation.
@@ 2 Preliminaries
Vision Transformers. ViTs [15] have introduced a transformer backbone to vision tasks by patchifying the input and viewing an image as a sequence of patch tokens and have proved its effectiveness on large-scale image classification tasks. While ViTs adopt an isotropic architecture, some following works on vision transformers [37; 25; 19; 40] adopt a pyramid-like hierarchical architecture that gradually downsamples the feature. The pyramid architecture is proved highly effective in classification and other downstream tasks. Apart from architectural improvements, some other works [3; 41] focuses on improving the Feed-Forward Network module in transformers.
Vision transformers are also mainstream backbones for denoising models. IPT [6] introduces an isotropic transformer backbone for denoising and other low-level tasks. Some later works [23; 22; 9] follow the isotropic convention, but other denoising works [38; 42] shift to U-Net backbones as their design. The pioneering work of U-ViT [1] and DiT [28] introduces full-transformer backbones to diffusion as denoisers.
Recent Advancements in Diffusion Transformers. Following DiTs, some works investigate the training and diffusion [16; 27] strategies of Diffusion Transformers. Other works focus on the design of the DiT backbone. DiffiT [8; 18] introduces a new fusion method for conditions; FiT [26] and VisionLLaMA [10] strengthens DiT by introducing LLM tricks including RoPE2D [34] and SwishGLU. These transformer-based diffusion works agree on adopting isotropic architectures on latents, i.e. the latent feature space is not downsampled throughout the whole diffusion model. The authors of DiT [28] even regard the inductive bias of U-Net as “not crucial”.
U-Nets for Diffusion. From canonical works [20; 33; 13; 29], the design philosophy of U-Net [30] is generally accepted in diffusion. Specifically, Stable Diffusion [29] uses a U-Net-based denoiser on the compressed latent space for high-resolution image synthesis, which is highly successful in manifold generative tasks. Some previous trials on diffusion transformers [4; 18; 11; 21] also adopt U-Net on pixel-space generation tasks; but strangely, they shifted to isotropic DiT-like structures for latent-space diffusion. Despite its popularity in pixel-space diffusion, the U-Net architecture is not widely accepted in recent transformer-oriented works on latent-space diffusion.
Motivated by this, we are dedicated to investigating the potential of Transformer-backboned U-Net on latent-space diffusion. It is noteworthy that our goal is significantly different from U-ViT [1]: U-ViT is an isotropic transformer architecture with shortcuts, but our work resort to true U-Net architectures that involves multiple stages of feature-map downsampling and upsampling.
@@ 3 Investigating U-Net DiTs in Latent
In fact, attention to downsampled tokens is not new. Previous works regarding vision transformers [17; 44] have proposed methods to downsample key-value pairs for computation cost reduction. Recent work on acceleration of diffusion models [32; 7] also applies key-value downsampling on Stable Diffusion models. But these works maintain the number of queries, and thus the downsampling operation is not completely performed. Besides, these downsampling measures usually involves a reduction of tensor size, which could result in a significant loss in information.
@@
Complexity analysis. Apart from the performance benefits, we are aware that adopting downsampled self-attention in the U-Shaped DiT could save as much as 1/3 of the model’s overall computation cost. We conduct a brief computation complexity analysis on the self-attention mechanism to explain where the savings come from.
@@
In the proposed self-attention on downsampled tokens, four sets of downsampled query-key-value tuples $4\times\left(Q_{\downarrow 2},K_{\downarrow 2},V_{\downarrow 2}\right)\in\mathbb{R}^{(\frac{N}{2})^{2}\times d}$ performs self-attention respectively. While each self-attention operation costs only 1/16 of full-scale self-attention, the total cost for downsampled self-attention is 1/4 of full-scale self-attention. 3/4 of the self-attention computation is saved via token downsampling.
@@ 4 Scaling the Model Up
Settings. We adopt the training setting of DiT. The same VAE (i.e. sd-vae-ft-ema) for latent diffusion models [29] and the AdamW optimizer is adopted. The training hyperparameters are kept unchanged, including global batch size 256, learning rate $41e-$ 4, weight decay 0, and global seed 0. The training is conducted with the training set of ImageNet 2012 [12]. We used 8 NVIDIA A100s (80G) to train U-DiT-B and U-DiT-L models. The training overhead is listed in the appendix.
Apart from the self-attention on downsampling as introduced in the toy experiment (Section 3), we further introduce a series of modifications to U-DiTs, including cosine similarity attention [24; 22], RoPE2D [34; 26; 10], depthwise conv FFN [38; 3; 44], and re-parametrization [14; 35]. The contribution of each modification is quantitatively evaluated in Sec. 9.
@@ 4.1 U-DiT at Larger Scales
Table 2: Comparing U-DiTs against DiTs on ImageNet 256$\times$256 generation. Experiments with supermarks are replicated according to the official code of DiT. We compare models trained for 400K iterations with the standard training hyperparameters of DiT. The performance of U-DiTs is outstanding: U-DiT-B could beat DiT-XL/2 with only 1/6 of inference FLOPs; U-DiT-L could outcompete DiT-XL/2 by $10$ FIDs.
ImageNet 256$\times$256
[…]
U-ViT-L [1] 76.4 21.22 6.10 67.64 0.615 0.633
U-ViT-XL [1] 113.0 18.35 5.75 76.59 0.632 0.630
DiT-XL/2 [28] 118.7 20.05 6.25 66.74 0.632 0.629
PixArt-$\alpha$-XL/2 [8] 118.4 24.75 6.08 52.24 0.612 0.613
DiffiT-XL/2 [18] 118.5 36.86 6.53 35.39 0.540 0.613
U-DiT-B (Ours) 22.2 16.64 6.33 85.15 0.642 0.639
U-DiT-L (Ours) 85.0 10.08 5.21 112.44 0.702 0.631
Table 3: Comparing U-DiTs against competitive diffusion architectures on ImageNet 256$\times$256 generation. Since different architectures use different training settings, we align them under the official 400K-iteration setting of DiT for a fair comparison. The proposed U-DiT series could outperform these models by large margins at fewer FLOPs. Experiments with supermarks include necessary modifications of the original work (detailed in the appendix).
@@
It could be concluded from Tab. 2 that all U-DiT models could outcompete their isotropic counterparts by considerable margins. Specifically, U-DiT-S and U-DiT-B could outperform DiTs of comparable size by $\sim 30$ FIDs; U-DiT-L could outperform DiT-XL/2 by $\sim 10$ FIDs. It is shocking that U-DiT-B could outcompete DiT-XL/2 with only 1/6 of the computation costs. In Tab. 3, we further demonstrate the advantage of U-DiTs over several competitive diffusion transformers [1; 28; 8; 18]. To present the advantage of our method better, we also include the performance of U-DiTs in an FID-50K versus FLOPs plot (Fig. 2). Apart from DiTs and U-DiTs, we also include other state-of-the-art methods: SiT [27] that proposes an interpolant framework for DiTs, and SiT-LLaMA [10] that combines state-of-the-art DiT backbone VisionLLaMA and SiT. The advantages of U-DiTs over other baselines are prominent in the plot. The results highlight the extraordinary scalability of the proposed U-DiT models.
@@ 4.2 Ablations
Next, we inspect other modifications that further refine U-DiTs (lower part of Tab. 9). Swin Transformer V2 [24] proposes a stronger variant of self-attention: instead of directly multiplying Q and K matrices, cosine similarities between queries and keys are used. We apply the design to our self-attention, which yields $\sim 2.5$FIDs of improvement. RoPE [34] is a powerful positional embedding method, which has been widely applied in Large Language Models. Following the latest diffusion transformer works [26; 10], we inject 2-dimensional RoPE (RoPE2D) into queries and keys right before self-attention. The introduction of RoPE2D improves performance by $\sim 2.5$FIDs. Some recent transformer works strengthen MLP by inserting a depthwise convolution layer between two linear mappings [38; 3; 44]. As the measure is proved effective in these works, we borrow it to our U-DiT model, improving $\sim 5$FIDs. As re-parametrization during training [14] could improve model performance, we apply the trick to FFN [35] and bring an additional improvement of $\sim 3.5$FIDs. Above all, based on these components, the proposed U-DiTs are further improved.